OpenStudy (anonymous):

In physics, Albert Einstein's famous formula for mass-energy equivalence defines an object's energy in terms of its mass and the speed of light. The formula is commonly written as E=MC 2 where E = energy, M = mass, and C = the speed of light. Rearrange the formula to solve for the speed of light (C).

OpenStudy (anonymous):

well u have e=mc right

OpenStudy (anonymous):

u want c alone but m is still with c

OpenStudy (anonymous):

soo u would kick m out to get c alone :(

OpenStudy (anonymous):

when m leaves c it goes under e cause e dont want any body to be with

OpenStudy (anonymous):

soo what do we have now ?

OpenStudy (wolf1728):

E = mc² we want to solve for speed of light so c² = E/m c = square root (E/m) (Here's a good E=mc² calculator) http://www.1728.org/einstein.htm